# Generalized Finite-time Optimal Control Framework in Stochastic Thermodynamics

### v2

@@ 2 Slow Driving Optimal Control
We consider a GFTOC problem of slow driving of the effective Lagrangian $\mathcal{L}^{*}$ in eq.˜4 from the initial $A_{\alpha}^{i}$ to the final control parameter $A_{\alpha}^{f}$ in a time $\tau$ that minimizes the EPR required for the driving process, which is implemented through external parametric control of $A_{\alpha}$. Here, the control of transition affinity $A_{\alpha}$ is generally used to denote multiple linearly independent control parameters that can implement this. Three linearly independent cases of affinity control correspond to independently contributed EPR, namely $\partial_{t}\psi_{E},\langle\dot{\Sigma}_{E}^{ex}\rangle$, and $\langle\dot{\Sigma}^{hk}\rangle$, quantified by the corresponding effective Lagrangians and relevant controllable affinities are $\{\lambda_{E}\}$, $\{-\log{\left(\rho_{i}/\rho_{i}^{E}\right)}\}$, and $\{F_{\gamma}\}$, respectively.
@@ 2.1 Geodesic structure
We consider slow driving of the Lagrangian, which physically implies that the system dynamics and statistics quickly adapt to the instantaneous control parameters. Hence, the slow driving formulation is valid under a timescale separation, precisely, only if the driving timescale is slower than the coupling timescale between the environment and the system or the relaxation timescale of the system, which quantifies how quickly the system adapt to the instantaneously imposed control parameters. Under the assumption of slow driving, we expand the Lagrangian to second-order terms in the rate of driving $\dot{A}_{\alpha}$. The driving Lagrangian reads [13, 14, 15]:
@@
such that the total driving EP is $\mathcal{S}_{drv}^{qs}=\Sigma_{drv}^{qs}=\int_{0}^{\tau}dt\>\mathcal{L}_{drv}^{*}$. Equation˜5 quantifies the driving kinetic energy, where the mass is given by the instantaneous curvature of the Lagrangian ($\partial_{{\alpha}}^{2}\mathcal{L}^{*}$). We adopt the shorthand notation $\partial_{{\alpha}}^{2}\mathcal{L}^{*}=\partial_{A_{\alpha}}^{2}\mathcal{L}^{*}=D_{\alpha}\left[2\cosh{(A_{\alpha}/2)}+\frac{1}{2}A_{\alpha}\sinh{(A_{\alpha}/2)}\right]=T_{\alpha}+\frac{1}{4}\mathcal{L}^{*}$.
The fEQ driven systems (with a larger $A_{\alpha}$) exhibit a higher mass, attributed to larger fluctuations ($T_{\alpha}$) and dissipation ($\mathcal{L}^{*}$). Physically, this signifies a higher resistance to the driving for fEQ systems, fig.˜1(a). fEQ driving is slowed due to increased resistance, with a singularity in the $A_{\alpha}\to\infty$ limit caused by divergence of the EPR and traffic. Importantly, the local curvature $\partial_{{\alpha}}^{2}\mathcal{L}^{*}$ generalizes the stochastic Fisher information for any $\mathcal{L}^{*}$. Choosing excess EP as the total dissipation cost shows the equivalence between mass and stochastic Fisher information for each state, discussed subsequently below. We have omitted the potential energy boundary term, namely, the instantaneous EPR, which does not depend on driving speed and time, and corresponds to the unavoidable instantaneous minimum EPR cost. This reduces the slow-driving optimization problem to a minimization problem of the driving kinetic energy (driving EPR). Its solution is the geodesic equation for $A_{\alpha}$,
@@
eq.˜6 generalizes the geodesic equation for the optimal control of fEQ systems that minimizes driving EPR [17]. It equivalently implies $\dot{A_{\alpha}}\sqrt{\partial_{{\alpha}}^{2}\mathcal{L}^{*}}={v}_{qs}$, where $v_{qs}$ is the quasi-static driving speed along the geodesic in the driving time $\tau$. The minimum action solution for optimal driving is called the geodesic and is denoted by $\mathcal{G}(A_{\alpha})$. It encodes the minimum distance path in the control parameter space, and its analytical form plays a crucial role. By definition, the geodesic implies $\mathcal{G}(A_{\alpha})=\arg\inf_{A_{\alpha}}\left(\Sigma_{drv}^{qs}\right)$, and the corresponding minimum EP is $\Sigma_{qs}^{*}=\inf_{A_{\alpha}}\left(\Sigma_{drv}^{qs}\right)$, under the initial and final boundary conditions $A_{\alpha}^{i}$ and $A_{\alpha}^{f}$. Thus, $\dot{A_{\alpha}}\propto 1/\sqrt{\partial_{{\alpha}}^{2}\mathcal{L}^{*}}$ implies that $A_{\alpha}$ has to be driven slower the further the system is from equilibrium. Physically, this means that for the validity of the optimal driving, each infinitesimally small driving timestep $\Delta t$ contributes equally to $\Sigma_{drv}^{qs}$.
@@
$\begin{split}\mathcal{G}(A_{\alpha})=\left(1-\frac{t}{\tau}\right)\mathcal{G}(A_{\alpha}^{i})+\frac{t}{\tau}\mathcal{G}(A_{\alpha}^{f})\end{split}$ (7)
@@
Figure 1: (a) The curvature of the Lagrangian $\partial_{\alpha}^{2}\mathcal{L}^{*}$ (mass) increases exponentially for the higher value of driving affinity $A_{\alpha}$. (b) $\mathcal{G}(F_{\alpha}):F_{\alpha}\to t$. Comparison between $\mathcal{G}^{cEQ}(F_{\alpha})$, $\mathcal{G}^{fEQ}(F_{\alpha})$ and $\mathcal{G}^{lin}(F_{\alpha})$ given by eq.˜9. For the fixed quasi-static driving speed $v_{qs}$, $F_{\alpha}^{f}-F_{\alpha}^{i}=1$ is considered for the cEQ $F_{\alpha}^{i}=1$ and fEQ $F_{\alpha}^{i}=3.5$ and the corresponding $\Delta\tau_{c}$ and $\Delta\tau_{f}$ are plotted; it is a combined pictorial visualization of eqs.˜9 and 13. Due to eq.˜13, the speed limit for the quasi-static slow-driving increases for fEQ systems, an effect attributed to the higher mass required for the driving of fEQ systems.
This simplification allows us to visualize the physical meaning of geodesic space. In particular, the optimal control problem for $A_{\alpha}$ with a varying instantaneous local mass is converted into one with a unit mass using the geodesic map $\mathcal{G}(A_{\alpha})$. Similarly, time is scaled in units of the driving time $\tau$. This mapping converts the non-quadratic optimal control problem in $A_{\alpha}$ space to a quadratic optimal control problem in $\mathcal{G}(A_{\alpha})$ with a scaled unit mass and a scaled unit time. This is a fundamental non-equilibrium scale-invariance relation for the optimal quasi-static driving of fEQ systems, and it delineates the underlying fundamental universality of all quasi-statically driven systems.
@@ 2.2 Exact analytical expressions for geodesics
In this section, we compute the exact analytical expressions for the geodesic for the control of transition affinity $A_{\alpha}$ for three different cases. We fix the amplitude of mass to 1 in the $A_{\alpha}\to 0$ limit. The first case is the linear geodesic $\mathcal{G}^{lin}$, which is computed using the equilibrium approximation of mass $\partial_{{A_{\alpha}}}^{2}\mathcal{L}^{*}=1$. The second case is the close-to-equilibrium geodesic $\mathcal{G}^{cEQ}$, which is computed using the close-to-equilibrium approximation. Up to $O(A_{\alpha}^{2})$ terms of $\partial_{{A_{\alpha}}}^{2}\mathcal{L}^{*}$, this implies $\partial_{{A_{\alpha}}}^{2}\mathcal{L}^{*}=1+A_{\alpha}^{2}/4$. The third case is the fEQ geodesic $\mathcal{G}^{fEQ}$, which is computed using the fEQ approximation of $\partial_{{A_{\alpha}}}^{2}\mathcal{L}^{*}$ using $\cosh{(a)}=\sinh{(a)}=\frac{1}{2}e^{a}$, implying $\partial_{{A_{\alpha}}}^{2}\mathcal{L}^{*}=\frac{1}{4}(A_{\alpha}+4)e^{\frac{1}{2}A_{\alpha}}$.
@@
In $A_{\alpha}\to 0$ limit, $\mathcal{G}^{lin}(A_{\alpha}),\mathcal{G}^{cEQ}(A_{\alpha})$, and $\mathcal{G}^{fEQ}(A_{\alpha})$ converge. However, the further the system is from equilibrium, the greater the quantitative difference is observed. The physical implication is a tighter bound on the exact optimal driving EP in eq.˜8, and the deviation of the optimal driving protocol eq.˜7 from a trivial linear interpolation between the initial and final control parameters. This highlights the importance of the exact fEQ geodesics.
@@ 2.3 Far-from-equilibrium generalizations of
Optimal mass transport theory
The geodesic hierarchy eq.˜10 implies the hierarchy $\left(\mathcal{G}^{fEQ}(A_{\alpha}^{f})-\mathcal{G}^{fEQ}(A_{\alpha}^{i})\right)^{2}\geq\left(\mathcal{G}^{cEQ}(A_{\alpha}^{f})-\mathcal{G}^{cEQ}(A_{\alpha}^{i})\right)^{2}\geq\left(A_{\alpha}^{f}-A_{\alpha}^{i}\right)^{2}$. Using eq.˜8, for a fixed $\tau$, a hierarchy on the driving EP reads,
@@
$\mathcal{G}^{lin}(A_{\alpha})$ is the special case that has been utilized in the OMTT [31, 32], and applied to ST to obtain $\Sigma_{qs}^{*lin}$ as the driving cost of EP, which is quadratic function of the change in the driving affinity [35, 43]. However, the computation of $\mathcal{G}$ takes into account the non-quadratic nature of the dissipation function $\mathcal{L}^{*}$, which results in the tightest and exact bound using eq.˜11. Moreover, the right-hand side of eq.˜8 is equal to the square of the $\mathcal{W}_{2}$ Wasserstein distance defined in the geodesic space of control parameters. Thus, eq.˜8 leads to equality between the driving EP in stochastic thermodynamics and the Wasserstein distance in OMTT,
$\Sigma_{qs}^{*}=\frac{\mathcal{W}_{2}^{2}}{2\tau}$ (12)
Therefore, our optimal control formulation goes beyond the existing quadratic counterparts formulated using the OMTT [35, 43, 44, 51]. Importantly, our slow-driving formulation also generalizes the connection between ‘thermodynamic geometry’ and ‘OMTT’ beyond known quadratic formulations. This extends the applicability of ‘OMTT’ to non-Euclidean (Riemannian) control parameter manifolds.
2.4 Speed limit for slow driving
The validity of slow driving is a key assumption in the optimal control framework developed so far. Equation˜8 reveals an inherent timescale associated with the driving process. Inverting eq.˜8, we define the quasi-static driving timescale:
@@ 2.4 Speed limit for slow driving
$\tau_{qs}^{*}$ is a quantitative measure of the time required for the given initial and final control parameters ($A_{\alpha}^{i}$ and $A_{\alpha}^{f}$) and the driving EP $\Sigma_{qs}^{*}$. It quantifies the timescale for the violation of the slow-driving assumption. Physically, it is equal to the square of the distance travelled along the geodesic in the control parameter space divided by the quasi-static EP supplied for the driving process. The quasi-static driving time reveals the fundamental tradeoff between the driving time and dissipation for the optimally driven process, and defines the speed limit for the slow-driving process. Importantly, the slow-driving assumption is violated for any $\tau$ and does not necessarily require $\tau\leq\tau_{qs}^{*}$. For the fixed value $\Delta A_{\alpha}=A_{\alpha}^{f}-A_{\alpha}^{i}$ and a fixed available budget for driving EP $\Sigma_{qs}^{*}$, eq.˜10 implies $\tau_{qs}^{*fEQ}>\tau_{qs}^{*cEQ}>\tau_{qs}^{*lin}$, see fig.˜1(b). Hence, fEQ systems require a larger quasi-static driving time, attributed to critical slowing due to higher traffic and EPR. This implies that fEQ systems are more prone to breaking the assumption of slow-driving, with prominent consequences discussed subsequently in section˜3.

### v3

@@ 2 Slow Driving Optimal Control
We consider a GFTOC problem of slow driving of the effective Lagrangian $\mathcal{L}^{*}$ in eq.˜4 from the initial $A_{\alpha}^{i}$ to the final control parameter $A_{\alpha}^{f}$ in a time $\tau$ that minimizes the EPR required for the driving process, which is implemented through external parametric control of $A_{\alpha}$. Here, the control of transition affinity $A_{\alpha}$ is generally used to denote multiple linearly independent control parameters that can implement this. Three linearly independent cases of affinity control correspond to independently contributed EPR, namely $\partial_{t}\psi_{E},\langle\dot{\Sigma}_{E}^{ex}\rangle$, and $\langle\dot{\Sigma}^{hk}\rangle$, quantified by the corresponding effective Lagrangians and relevant controllable affinities are $\{\lambda_{E}\}$, $\{-\log{\left(\rho_{i}/\rho_{i}^{E}\right)}\}$, and $\{F_{\gamma}\}$, respectively, see eq.˜2.
@@ 2.1 Geodesic structure
We consider slow driving of the Lagrangian, which physically implies that the system dynamics and statistics quickly adapt to the instantaneous control parameters. Hence, the slow driving formulation is valid under a timescale separation, precisely, only if the driving timescale is slower than the coupling timescale between the environment and the system or the relaxation timescale of the system, which quantifies how quickly the system adapts to the instantaneously imposed control parameters. Under the assumption of slow driving, we expand the Lagrangian to second-order terms in the rate of driving $\dot{A}_{\alpha}$. The driving Lagrangian reads [20, 21, 22]:
@@
such that the total driving EP is $\mathcal{S}_{drv}^{qs}=\Sigma_{drv}^{qs}=\int_{0}^{\tau}dt\>\mathcal{L}_{drv}^{*}$.
Equation˜5 quantifies the driving kinetic energy, where the mass is given by the instantaneous ‘local curvature’ of the Lagrangian ($\partial_{{\alpha}}^{2}\mathcal{L}^{*}$). We adopt the shorthand notation $\partial_{{\alpha}}^{2}\mathcal{L}^{*}=\partial_{A_{\alpha}}^{2}\mathcal{L}^{*}$, which satisfies
$\partial_{{\alpha}}^{2}\mathcal{L}^{*}=D_{\alpha}\left[2\cosh{\left(\frac{A_{\alpha}}{2}\right)}+\frac{1}{2}A_{\alpha}\sinh{\left(\frac{A_{\alpha}}{2}\right)}\right]=T_{\alpha}+\frac{1}{4}\mathcal{L}^{*}$ (6)
The fEQ driven systems (with a larger $A_{\alpha}$) exhibit a higher mass, attributed to larger fluctuations ($T_{\alpha}$) and dissipation ($\mathcal{L}^{*}$). Physically, this signifies a higher resistance to the driving for fEQ systems, fig.˜1(a) plots eq.˜6. The fEQ driving is critically slowed due to increased resistance, with a singularity in the $A_{\alpha}\to\infty$ limit, which is caused by divergence of the EPR and traffic. Importantly, the local curvature $\partial_{{\alpha}}^{2}\mathcal{L}^{*}$ generalizes the stochastic Fisher information for any $\mathcal{L}^{*}$, thereby going beyond the existing formulation in information geometry [38]. Choosing excess EP as the total dissipation cost shows the equivalence between mass and stochastic Fisher information for each state, which is the predominantly studied dissipation function in information geometry [38], discussed subsequently below in section˜5.3. We have omitted the potential energy boundary term, namely, the instantaneous EPR, which does not depend on driving speed and time, and corresponds to the unavoidable instantaneous minimum EPR cost, in the limit $\tau\to\infty$. This reduces the slow-driving optimization problem to a minimization problem of the driving kinetic energy (driving EPR). Its solution is the geodesic equation for $A_{\alpha}$,
@@
eq.˜7 generalizes the geodesic equation for the optimal control of fEQ systems that minimizes driving EPR [24]. It equivalently implies $\dot{A_{\alpha}}\sqrt{\partial_{{\alpha}}^{2}\mathcal{L}^{*}}={v}_{qs}$, where $v_{qs}$ is the quasi-static driving speed along the geodesic in the driving time $\tau$. The minimum action solution for optimal driving is called the geodesic and is denoted by $\mathcal{G}(A_{\alpha})$. It encodes the minimum distance path in the control parameter space, and its analytical form plays a crucial role. By definition, the geodesic implies $\mathcal{G}(A_{\alpha})=\arg\inf_{A_{\alpha}}\left(\Sigma_{drv}^{qs}\right)$, and the corresponding minimum EP is $\Sigma_{qs}^{*}=\inf_{A_{\alpha}}\left(\Sigma_{drv}^{qs}\right)$, under the initial and final boundary conditions $A_{\alpha}^{i}$ and $A_{\alpha}^{f}$. Thus, $\dot{A_{\alpha}}\propto 1/\sqrt{\partial_{{\alpha}}^{2}\mathcal{L}^{*}}$ implies that $A_{\alpha}$ has to be driven slower the further the system is from equilibrium. Physically, this means that for the validity of the optimal driving, each infinitesimally small driving timestep $\Delta t$ contributes equally to $\Sigma_{drv}^{qs}$.
@@
$\begin{split}\mathcal{G}(A_{\alpha})=\left(1-\frac{t}{\tau}\right)\mathcal{G}(A_{\alpha}^{i})+\frac{t}{\tau}\mathcal{G}(A_{\alpha}^{f}).\end{split}$ (8)
@@
Figure 1: (a) The curvature of the Lagrangian $\partial_{\alpha}^{2}\mathcal{L}^{*}$ (mass given by eq.˜6) increases exponentially for the higher value of driving affinity $A_{\alpha}$. (b) $\mathcal{G}(A_{\alpha}):A_{\alpha}\to t$. Comparison between $\mathcal{G}^{cEQ}(A_{\alpha})$, $\mathcal{G}^{fEQ}(A_{\alpha})$ and $\mathcal{G}^{lin}(A_{\alpha})$ given by eq.˜10. For the fixed quasi-static driving speed $v_{qs}$, $A_{\alpha}^{f}-A_{\alpha}^{i}=1$ is considered for the cEQ $A_{\alpha}^{i}=1$ and fEQ $A_{\alpha}^{i}=3.5$ and the corresponding $\Delta\tau_{c}$ and $\Delta\tau_{f}$ are plotted; it is a combined pictorial visualization of eqs.˜10 and 15. Due to eq.˜15, the geometric speed limit for the quasi-static slow-driving increases for fEQ systems, an effect attributed to the higher mass required for the driving of fEQ systems.
This simplification allows us to visualize the physical meaning of geodesic space using eqs.˜8 and 9. In particular, the optimal control problem for $A_{\alpha}$ with a varying instantaneous local mass is converted into one with a unit mass using the geodesic map $\mathcal{G}(A_{\alpha})$. Similarly, time is scaled in units of the driving time $\tau$. This mapping converts the non-quadratic optimal control problem in $A_{\alpha}$ space to a quadratic optimal control problem in $\mathcal{G}(A_{\alpha})$ with a scaled unit mass and a scaled unit time. This is a fundamental non-equilibrium scale-invariance relation for the optimal quasi-static driving of fEQ systems, and it delineates the underlying fundamental universality of all quasi-statically driven systems.
@@ 2.2 Exact analytical expressions for geodesics
In this section, we compute the exact analytical expressions for the geodesic for the control of transition affinity $A_{\alpha}$ for three different cases. We fix the amplitude of mass to 1 in the $A_{\alpha}\to 0$ limit. The first case is the linear geodesic $\mathcal{G}^{lin}$, which is computed using the equilibrium approximation of mass $\partial_{{A_{\alpha}}}^{2}\mathcal{L}^{*}=1$. The second case is the close-to-equilibrium geodesic $\mathcal{G}^{cEQ}$, which is computed using the close-to-equilibrium approximation. Up to $O(A_{\alpha}^{2})$ terms of $\partial_{{A_{\alpha}}}^{2}\mathcal{L}^{*}$, this implies $\partial_{{A_{\alpha}}}^{2}\mathcal{L}^{*}=1+A_{\alpha}^{2}/4$. The third case is the fEQ geodesic $\mathcal{G}^{fEQ}$, which is computed using the fEQ approximation of $\partial_{{A_{\alpha}}}^{2}\mathcal{L}^{*}$ using $\cosh{(a)}\approx\sinh{(a)}\approx\frac{1}{2}e^{a}$, implying $\partial_{{A_{\alpha}}}^{2}\mathcal{L}^{*}=\frac{1}{4}(A_{\alpha}+4)e^{\frac{1}{2}A_{\alpha}}$.
@@
In $A_{\alpha}\to 0$ limit, $\mathcal{G}^{lin}(A_{\alpha}),\mathcal{G}^{cEQ}(A_{\alpha})$, and $\mathcal{G}^{fEQ}(A_{\alpha})$ converge. However, the further the system is from equilibrium, the greater the quantitative difference is observed. The physical implication is a tighter bound on the exact optimal driving EP in eq.˜9, and the deviation of the optimal driving protocol eq.˜8 from a trivial linear interpolation between the initial and final control parameters. This highlights the importance of the exact fEQ geodesics.
@@ 2.3 Far-from-equilibrium generalizations of
Optimal transport theory
The geodesic hierarchy eq.˜11 implies the hierarchy $\left(\mathcal{G}^{fEQ}(A_{\alpha}^{f})-\mathcal{G}^{fEQ}(A_{\alpha}^{i})\right)^{2}\geq\left(\mathcal{G}^{cEQ}(A_{\alpha}^{f})-\mathcal{G}^{cEQ}(A_{\alpha}^{i})\right)^{2}\geq\left(A_{\alpha}^{f}-A_{\alpha}^{i}\right)^{2}$, see fig.˜1(b). For a fixed $\tau$, a hierarchy on the driving EP obtained using eq.˜9 reads,
@@
$\mathcal{G}^{lin}(A_{\alpha})$ is the special case that has been utilized in optimal transport theory [39, 40], and applied to ST to obtain $\Sigma_{qs}^{*lin}$ as the driving cost of EP, which is quadratic function of the change in the driving affinity [43, 53]. In this aspect, optimal transport theory is a cEQ ‘hybrid’ case that deals with a ‘state-space’ representation, for which the mass for the excess EP is utilized (see section˜5.3)  [^5] There is a minor technical difference that is not of fundamental importance, since the mass defined for a control parameter is the inverse of the mass defined for its dual conjugate; see section 5.3., but a quadratic dependence on the driving-affinity incorporates the housekeeping EPR (see section˜5.4).
However, the computation of $\mathcal{G}$ takes into account the non-quadratic nature of the dissipation function $\mathcal{L}^{*}$, which results in the tightest and exact bound using eq.˜12. Moreover, the right-hand side of eq.˜9 is equal to the square of the $\mathcal{W}_{2}$ Wasserstein distance defined as,
$\begin{split}\mathcal{W}_{2}=\mathcal{G}(A_{\alpha}^{f})-\mathcal{G}(A_{\alpha}^{i}),\end{split}$ (13)
the distance in the geodesic space of the control parameters. This is followed by the definition of $\mathcal{W}_{2}$ and recalling the definition of the geodesic, which physically corresponds to the minimization of distance, and is mathematically equal to the $L^{2}-{norm}$. The definition eq.˜13 of $\mathcal{W}_{2}$ itself is novel, and accounts for the geometric curvature for any control parameter. The existing cEQ optimal transport theory methodology focuses on the control of the state-space. Thus, eq.˜9 leads to equality between the driving EP in ST and the Wasserstein distance in OT,
$\Sigma_{qs}^{*}=\frac{\mathcal{W}_{2}^{2}}{2\tau}.$ (14)
Therefore, our optimal control formulation goes beyond the existing quadratic counterparts formulated using the optimal transport theory [43, 53, 54, 58]. Importantly, our slow-driving formulation also reveals the connection between ‘thermodynamic geometry’ and ‘optimal transport theory’ beyond known quadratic formulations. This extends the applicability of ‘optimal transport theory’ to any non-Euclidean (Riemannian) control parameter manifolds.
2.4 Geometric speed limit for slow driving:
a time-dissipation-geometry tradeoff relation
The validity of slow driving is a key assumption in the optimal control framework developed so far. Equation˜9 reveals an inherent timescale associated with the driving process. Inverting eq.˜9, we define the quasi-static driving timescale:
@@ 2.4 Geometric speed limit for slow driving:
$\tau_{qs}^{*}$ is a quantitative measure of the time required for the given initial and final control parameters ($A_{\alpha}^{i}$ and $A_{\alpha}^{f}$) and the driving EP $\Sigma_{qs}^{*}$. It quantifies the timescale for the violation of the slow-driving assumption. Physically, it is equal to the square of the distance travelled along the geodesic in the control parameter space divided by the quasi-static EP supplied for the driving process.
The quasi-static driving time reveals the fundamental tradeoff between the driving time, dissipation (EP) and geometry for the optimally driven process, and defines the geometric speed limit for the slow-driving process. Therefore, eq.˜15 is a fundamental time-dissipation-geometry tradeoff relation. Importantly, the slow-driving assumption is violated for any $\tau$ and does not necessarily require $\tau\leq\tau_{qs}^{*}$. For the fixed value $\Delta A_{\alpha}=A_{\alpha}^{f}-A_{\alpha}^{i}$ and a fixed available budget for driving EP $\Sigma_{qs}^{*}$, eq.˜11 implies,
$\tau_{qs}^{*fEQ}>\tau_{qs}^{*cEQ}>\tau_{qs}^{*lin},$ (16)
see fig.˜1(b). Hence, fEQ systems require a larger quasi-static driving time, attributed to critical slowing due to higher mass. This implies that fEQ systems are more prone to breaking the assumption of slow-driving, with prominent consequences discussed subsequently in section˜3.